গাণিতিক যুক্তি শিক্ষক নিবন্ধন +৪ more ১৮ তম শিক্ষক নিবন্ধন একটি বর্গক্ষেত্রের এক বাহুর দৈর্ঘ্যের দ্বিগুণ 6 একক হলে, এর কর্ণের দৈর্ঘ্যের দ্বিগুণ কত একক? ক 3 1 32 2 62 3 82 সঠিক উত্তর 2 (62)
Bank Preparation Bank English +৮ more Bangladesh Bank Assistant General Director- 2018 We were at a disadvantage_______that we did not had a very good knowledge of the language the others were using- ক by 1 in 2 with 3 for সঠিক উত্তর 1 (in)
Bank Preparation Bank English +৫ more Bangladesh Bank Assistant General Director- 2018 PyongYang and Washington have finally come to a consensus.'this is an example of- ক metaphor 1 metonym 2 euphemism 3 dysphemism সঠিক উত্তর 1 (metonym)
Bank Preparation Bank English +৫ more Bangladesh Bank Assistant General Director- 2018 Translet to English: তার কোনো জোরালো রাজনৈতিক আদর্শ নেই- ক He has no political axe to grind 1 He has no political will to move on. 2 He has no political idol to follow 3 He has no strong political stand to speak up সঠিক উত্তর ক (He has no political axe to grind)
Bank Preparation Bank Math +১ more A coin is tossed four times. What is the probability of getting head on all tosses? ক 1/4 1 1/8 2 1/16 3 3/16 সঠিক উত্তর 2 (1/16) ব্যাখ্যা: Probability of getting head on all tosses = 1/2 × 1/2 × 1/2 × 1/2 = 1/16
Bank Preparation Bank English +৫ more Bangladesh Bank Assistant General Director- 2018 If Asylum: Shelter then- ক Hospice:Exile 1 Harbour:Concealment 2 Stronghold:Defence 3 Palisade:Display সঠিক উত্তর 2 (Stronghold:Defence)
Bank Preparation Bank English +৭ more Bangladesh Bank Assistant General Director- 2018 Yhe antonym of 'inspid' is- ক brave 1 exciting 2 cold 3 dull সঠিক উত্তর 1 (exciting)
Bank Preparation Bank Math +১ more A fair coin is flipped three times. What is the probability that the coin lands head each time? ক 1/2 1 1/3 2 1/4 3 1/8 সঠিক উত্তর 3 (1/8) ব্যাখ্যা: All possible outcome = {HHH, HHT, HTT, HTH, THH, TTH, THT, TTT} = 8 It will be head every time, this occurs 1 time ∴ Probability = 1/8
Bank Preparation Bank English +৭ more Bangladesh Bank Assistant General Director- 2018 The synonym of 'sanguine' is- ক optimistic 1 restless 2 hopeless 3 bloody সঠিক উত্তর ক (optimistic)
Bank Preparation Bank Math +১ more What is the probability of impossible events? ক 1 1 0 2 >1 3 <1 সঠিক উত্তর 1 ব্যাখ্যা: Impossible events can't occur. The probability of an impossible event is 0.
Bank Preparation Bank English +৫ more Bangladesh Bank Assistant General Director- 2018 'Bravity is the soul of wit' এর যথাযথ বাংলা অনুবাদ ? ক অতি অল্প হইল 1 মানিকের খানিক ভাল 2 প্রাণের কথা বুকে বাজে 3 কথা কম কাজ বেশি সঠিক উত্তর 1 (মানিকের খানিক ভাল)
গাণিতিক যুক্তি শিক্ষক নিবন্ধন +১ more ১৮ তম শিক্ষক নিবন্ধন একটি খাড়া খুঁটি মাটি থেকে 5 মিটার উপরে ভেঙে বিচ্ছিন্ন না হয়ে অন্যপ্রান্ত ভূমিতে 5 মিটার দূরত্বে স্পর্শ করলে, খুঁটিটির উচ্চতা কত? ক 25 1 10 2 5(1+2) 3 5 সঠিক উত্তর 2 (5(1+2) )
Bank Preparation Bank Math +১ more What is the probability that an integer selected at random from those between 20 and 100 inclusive is a multiple of 15? ক 5/79 1 2/27 2 5/81 3 6/79 সঠিক উত্তর 2 (5/81) ব্যাখ্যা: Multiple of 15 from 20 to 100 is 30, 45, 60, 75, 90 = 5 ∴ Probability = 5/81
Bank Preparation English Language +৫ more Bangladesh Bank Assistant General Director- 2018 Special drawing right (sdrs) is reletated to- ক IMF 1 IDB 2 ACD 3 HSBC সঠিক উত্তর ক (IMF)
গাণিতিক যুক্তি শিক্ষক নিবন্ধন +৬ more ১৮ তম শিক্ষক নিবন্ধন x5p=1 হলে, p এর মান কত ? ক 0 1 1 2 -5 3 5 সঠিক উত্তর ক
Bank Preparation Bank Math +১ more Jennifer flipped a coin three times and got heads each time. What is the probability that she gets heads on the next flip? ক 1 1 1/16 2 1/2 3 0 সঠিক উত্তর 2 (1/2) ব্যাখ্যা: Every time a coin flips, its independent outcome has a 50-50 chance. So, the probability is 1/2
Bank Preparation Bank Math +১ more Wendy has 5 pairs and 8 shirts. How many different combinations can she make with these items? ক 13 1 24 2 40 3 21 সঠিক উত্তর 2 (40) ব্যাখ্যা: Wendy can make = 5 × 8 = 40 different combinations.
English Language ব্যাংক নিয়োগ +৭ more Bangladesh Bank Assistant General Director- 2018 If you can win his attention _____ for you. ক the so much better 1 the better so much 2 so much better 3 so much for better সঠিক উত্তর 2 (so much better)
Bank Preparation Bank Math +১ more A certain jar contains 60 jellybeans 22 white, 18 green, 11 yellow, 5 red and 4 purple. If a jellybeans is to be chosen at random, what is the probability that the jellybean will be neither red nor purple ? ক 0.09 1 0.85 2 0.54 3 0.91 সঠিক উত্তর 1 (0.85) ব্যাখ্যা: Solution: Total = 60 White = 22 Green = 18 Yellow = 11 Red = 5 Purple = 4 লাল ও বেগুনি জেলি সংখ্যা = 5 + 4 = 9 লাল ও বেগুনি জেলি উঠার সম্ভাবনা = 9/60 লাল ও বেগুনি জেলি না উঠার সম্ভাবনা = 1 - 9/60 = (60 - 9)/60 = 51/60 = 0.85
গাণিতিক যুক্তি শিক্ষক নিবন্ধন +২ more ১৮ তম শিক্ষক নিবন্ধন একটি রেখাংশের উপর অঙ্কিত বর্গক্ষেত্র ঐ রেখাংশের এক-তৃতীয়াংশের উপর অঙ্কিত বর্গক্ষেত্রের কত গুণ? ক ৩ গুণ 1 ৬ গুণ 2 ৯ গুণ 3 দ্বিগুণ সঠিক উত্তর 2 (৯ গুণ)
গাণিতিক যুক্তি শিক্ষক নিবন্ধন +২ more ১৮ তম শিক্ষক নিবন্ধন দুটি ক্রমিক সংখ্যার বর্গের অন্তর 119 হলে, ছোট সংখ্যাটি কত? ক 38 1 59 2 60 3 78 সঠিক উত্তর 1 (59)
Bank Preparation Bank Math +১ more A boat covers 143 km upstream in 13 hours and the same distance downstream in 11 hours. What is the speed (in km/hr) of the boat in still (without stream) water? ক 10 km/hr 1 12 km/hr 2 14 km/hr 3 8 km/hr সঠিক উত্তর 1 (12 km/hr) ব্যাখ্যা: Solution: Let, the speed (in km/hr) of the boat in still (without stream) water is x km/hr and speed of stream is y km/hr Now 143/(x - y) = 13 ⇒ 13 (x - y) = 143 ⇒ x - y = 11..........(1) 143/(x + y) = 11 ⇒ 11 (x + y) = 143 ⇒ x + y = 13..........(2) (1) + (2) ⇒ x - y + x + y = 13 + 11 ⇒ 2x = 24 ∴ x = 12 km/hr
Bank Preparation Bank Math +১ more 10 minutes after a plane leaves the airport, it is reported to be 40 miles away. What is the average speed in miles per hour of the plane? ক 560 1 400 2 240 3 200 সঠিক উত্তর 2 (240) ব্যাখ্যা: Solution: গড় বেগ = {40/(10/60)} miles/hour = 40 × 6 miles/hour = 240 miles/hour
Bank Preparation Bank Math +১ more A person crosses a 600 m long street in 5 minutes. What is his speed in km per hour? ক 3.6 1 7.2 2 8.4 3 10 সঠিক উত্তর 1 (7.2) ব্যাখ্যা: Solution: 600 m = 600/1000 km = 3/5 km 5 minutes = 5/60 = 1/12 hr Speed = (3/5)/(1/12) = (3 × 12)/5 km/hr = 36/5 km/hr = 7.2 km/hr